OpenStudy (anonymous):

1 1 - -- ( 8 - ) 3 7 A. 5 - 2 -- 7 3 B. - - 8 1 C. - -- 21 D. 1 - 3

OpenStudy (anonymous):

What's the question ? \[-\frac{ 1 }{ 3 }\times 8\frac{ 1 }{ 7 }\] ?

OpenStudy (anonymous):

you times both of them. or divid it not sure

OpenStudy (anonymous):

The first thing we need to do is turn that mixed fraction into an improper fraction. The mixed fraction is \[8\frac{ 1 }{ 7 }\] Do you know how to do this?

OpenStudy (anonymous):

nope thats why im on here

OpenStudy (anonymous):

Okay lets think about it in terms of a 7 slice pizza. We have 8 full pizzas and one extra slice. We need to figure out how many slices we have in total. Your thoughts on this? Mathematically... to change a mixed fraction to an improper fraction we need to get a common denominator and add them. \[8\frac{ 1 }{ 7 }=8+\frac{ 1 }{ 7 }\] Let me know if this makes sense so far, then I'll explain further

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what exactly is confusing you. I don't think I can explain it much better.

OpenStudy (anonymous):

ill just guess

OpenStudy (anonymous):

Or you can actually try to apply yourself and learn. I will help

OpenStudy (anonymous):

If we have 8 pizzas and each one has 7 slices. How many total slices do we have?

OpenStudy (anonymous):

This in not an answer site. It is a tutor site.

OpenStudy (anonymous):

nice to know, did i ask for ur advice ? nope